OpenStudy (anonymous):

Find the integral: ∫√(6x-7) dx Please would like to know an explanation... I will appreciate, thanks!! ;)

OpenStudy (anonymous):

a simple u-substitution will do the trick: u = 6x-7, so du = 6dx try that....

OpenStudy (anonymous):

Please help me, to get into the answer.. I will appreciate. Thanks!

jimthompson5910 (jim_thompson5910):

use dpaInc's hint to go from ∫√(6x-7) dx to (1/6)*∫√u du

OpenStudy (anonymous):

when you substitute u and du into the equation you will go from \[\int\limits_{}^{} (\sqrt{6x-7}dx\] to \[\int\limits_{}^{}\sqrt{u} \frac{ du }{ 6 }\] because {u = 6x-7} and {du = 6dx} so divide both sides by 6 and you will be able to substitute for dx... (du/6) Once you understand that.. integrate the equation and you will get, \[\frac{ 2u ^{3/2} }{ 3 } (\frac{ 1 }{ 6 })\] Substitute u back into the equation you will get... \[\frac{ 2(6x-7)^{3/2} }{ 3 }(\frac{ 1 }{ 6 })\] Solved.

OpenStudy (anonymous):

@rocal2

OpenStudy (anonymous):

Sorry I forgot one part, add +C to the end of your answer for your constant after integrating.

OpenStudy (anonymous):

Thanks!! I like it, I will try another one....

OpenStudy (anonymous):

Your Welcome